OpenStudy (artfis):

The water output rate for the Jackson family's sprinkler was 15L per hour. The water output rate for the Simmons family's sprinkler was 30L per hour. The families used their sprinklers for a combined total of 55L hours, resulting in a total water output of 1425L. How long was each sprinkler used?

OpenStudy (artfis):

What would be the formula? That's all I need.

OpenStudy (anonymous):

Can you come up with some variables first?

OpenStudy (kropot72):

J = number of hours for Jackson's sprinkler S = number of hours for Simmon's sprinkler J + S = 55 Therefore S = 55 - J ......................(1) 15J + 30S = 1425 ..........................(2) Now just substitute the value of S from equation (1) into equation (2) and solve to find J.

OpenStudy (anonymous):

Figure out what you want your variables to represent.

OpenStudy (mathstudent55):

The Jackson family used x hours The Simmons family used y hours liters/hour times hours = liters used Water used: Jackson family: 15x Simmons family: 30y Total water used: 15x + 30y = 1425 Total time used: x + y = 55 Solve system of equations: 15x + 30y = 1425 x + y = 55

OpenStudy (artfis):

Thanks I see that I tried to solve for y first, and it got all jacked up.

OpenStudy (kropot72):

You're welcome :)